OpenStudy (anonymous):

SOME ONE PLEASE HELP y=1/2x-5 i have to graph it and there is only two points

OpenStudy (johnweldon1993):

Well you can see that this is a linear function.....just a y-intercept of -5 and an slope of 1/2.... so 1 point would be (0 , -5) ...gotten by plugging in 0 for 'x' y = 1/2(0) - 5 y = 0 - 5 y = -5 Now to get the other point you would need...just plug in 0 for 'y' this time and solve for x 0 = 1/2x - 5 5 = 1/2x 10 = x so your second point would be (10 , 0) So now you have 2 points....and you have a straight line function...so just use a straight line to connect these 2 points...and that will be your graph